OpenStudy (anonymous):

a sinusoidal electromagnetic wave with a frequency of 1.5x 10 ^ 15 Hz propagates in vacuum in the positive y-axis direction. Oscillating magnetic field parallel to the x axis with an amplitude of 3 mT. At y = 0 and t = 0, the large magnetic field is 0. Determine the function of the wave vector of the magnetic field and electric field at the time t.

OpenStudy (mos1635):

E=Eo*sin 2p(t/T-x/λ) B=Bo*sin 2p(t/T-x/λ) Eo/Bo=c

OpenStudy (anonymous):

How to get the Eo?

OpenStudy (mos1635):

Bo=3mT=3*10^-3 Tesla c=3*10^8

OpenStudy (anonymous):

What is relation Bo with Eo?

OpenStudy (mos1635):

Eo/Bo=c

OpenStudy (anonymous):

wow, thanks..

OpenStudy (mos1635):

wlc

OpenStudy (anonymous):

yep